OpenStudy (anonymous):

Is anyone here studying for Ap Calculus AB? Please Join us for a study group!

OpenStudy (anonymous):

I sure am

OpenStudy (anonymous):

Awesome Do you want to study together?

OpenStudy (anonymous):

sure but I gotta tell you I'm no good

OpenStudy (anonymous):

ok. I'm not that good either...but I think it will help to study with other people

OpenStudy (anonymous):

okeydoke

OpenStudy (anonymous):

You know I got ask....to better our conversation.....if you are a girl or a guy

OpenStudy (anonymous):

just to make it swift-ier

OpenStudy (anonymous):

I don't get it but I'm a gal

OpenStudy (anonymous):

oh hehe, I'm sorry..it's just me i guess. Anywho, girl here too....so let's start studying?

OpenStudy (anonymous):

sure watcha got?

OpenStudy (anonymous):

you know anything about volume of a revolution?

OpenStudy (anonymous):

i'm stuck on that....preferably volume of a cone

OpenStudy (anonymous):

really? I haven't done the volume of a cone, I'm on Application of integrals and integrals

OpenStudy (anonymous):

oh...the volume stuff is application of derivative....

OpenStudy (anonymous):

Volumes of the derivatives yeah that's in my application of integrals

OpenStudy (anonymous):

or volumes of revolutions*

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, so I have this question I'm stuck on...wait let me post it

OpenStudy (anonymous):

#2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Is it safe to assume that @S.W.'s answer is enough?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

you see i did look up that website and saw it

OpenStudy (anonymous):

I am only assuming it is correct, then that would answer part a

OpenStudy (anonymous):

but i didn't find it helpful because the question was different.. it said to compare it to geometric questions...however, the book says to use fundamental theorm.

OpenStudy (anonymous):

now...it could be my confusion with the topic, but i don't think these two are the same thing...

OpenStudy (anonymous):

so does anyone know what the fundamental theorem of calculus is?

OpenStudy (anonymous):

If \(F(x)=f'(x)\), then \(\displaystyle \int_a^b f(x)~dx=F(b)-F(a)\). What the question is asking of you is to find the volume using the disk method integration setup. Part (b) asks you to compare the answer to the volume you find by using the formula for the volume of a cone, \(V=\dfrac{1}{3}\pi r^2h\). The answer to part (c) is clear if you just compare the formula for cylinders and cones.

OpenStudy (anonymous):

in this case, the disk method is done vertically?